OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, -1) and perpendicular to y = 1/2x + 3.

OpenStudy (anonymous):

y = -2x - 7 y = -2x +5 y = 2x + 5 y = 2x - 7

OpenStudy (anonymous):

y = -2x - 7